THE EQUATION OF THE PARABOLA WHOSE FOCUS IS THE ORIGIN AND THE TANGENT AT THE VERTEX ISย  ย  ย  X-Y+1=0 IS

Grade:12th pass

1 Answers

Kailas Daud
6 years ago
x+y=0;x=-yย  /x-y+1=0 then 2x=-1,x=-1/2 or y=1/2 A=(-1/2,1/2) then z=(-1,1)ย  ย  ย SO,x-y+k=0ย  -1-1+k=0 k=+2ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  ย  (PM)2=(PN)2
ย  ย  ย  ย  ย  ย  X2+Y2+2XY-4X+4Y+4=0
ย